OpenStudy (anonymous):

Which expression gives the solutions of –5 + 2x² = –6x?

OpenStudy (anonymous):

2x^2 + 6x - 5 = 0

hero (hero):

Post the expressions.

OpenStudy (anonymous):

sorry thats not it

OpenStudy (anonymous):

Crap I messed up the formatting, the 4*1*5/2 isn't supposed to be an exponent :P

hero (hero):

pylauzier, I'm certain that that is not how the expression looks

OpenStudy (anonymous):

It would be - (4*2*(-5))

hero (hero):

\[\frac{-6 \pm \sqrt{6^2 - 4(2)(-5)}}{4}\]

OpenStudy (anonymous):

You realize that answer (minus the negative sign mistake) is identical to mine, right?

OpenStudy (anonymous):

Usually it's a good practice to simplify the ax^2 term to a=1.